# DocFusion: A Unified Framework for Document Parsing Tasks

### v1

@@ 5.1 Implementation Details
Model size OCR MER TR
[…]
Pix2tex (2022) - - - 76.5 41.7 95.9 - -
Texify (2023) 312M - - 88.6 71.7 97.8 - -
UniMERNet (2024a) 325M - - 99.0 89.5 99.7 - -
Qwen-VL-PLUS (2023) - 85.3 0.120 - - - - -
Qwen-VL-OCR (2023) - 94.9 0.055 - - - - -
StructEqTable (2024) 938M - - - - - 90.6 89.3
GOT (2024) 580M 86.7 0.115 87.7 67.3 97.8 86.9 81.6
DocFusion 289M 99.1 0.007 98.7 94.2 99.8 92.1 92.5
Table 2: Comparison of DocFusion with other models on three recognition tasks. Note: Due to differences in training styles across models, line break were consistently removed when calculating BLEU and Edit Distance.
[…]
DETR (2020) 41M 87.1 91.6 89.3 95.9 96.2 96.0 3.7
DocLayout-YOLO (2024c) 20M 86.7 91.1 88.9 94.4 95.5 95.0 85.2
DocFusion 289M 88.0 88.4 88.2 96.8 96.2 96.4 7.5
Table 3: The performance of the models on DLA, where DocLayNet-Scientific refers to the scientific document subset of the Doclaynet. Note: DETR and DocLayout-YOLO are limited to object detection tasks only.
5.2 Evaluation Metrics
5.2.1 Evaluation for Recognition
We employ traditional metrics such as BLEU Papineni et al. (2001) and Edit Distance Levenshtein (1966) to evaluate generated sequences. Additionally, we introduce task-specific metrics like CDM Wang et al. (2024b) and CSR to better assess the quality and usability of LaTeX-based outputs.
BLEU: The BLEU score is used for evaluating machine-generated text, measuring n-gram overlap with reference texts while incorporating a brevity penalty to ensure balanced outputs.
Edit distance: Also known as Levenshtein distance, measures the minimum number of operations insertions, deletions, or substitutions required to transform one string into another.
CSR: This score refers to the percentage of generated LaTeX outputs that can be successfully compiled into PDF. It reflects the correctness of the model’s predictions and practical usability.
ExpRate: The ExpRate Li et al. (2022) measures the proportion of samples where the predicted text matches the reference text without any errors.
CDM: The CDM evaluates MER by comparing image-rendered expression at the character level with spatial localization, ensuring fairness and accuracy over text-based metrics like BLEU.
5.2.2 Evaluation for Detection
Since the DLA task in DocFusion does not use confidence scores, we did not use the widely adopted Average Precision (AP) metric from the object detection field. Instead, we focus on the following metrics:
Precision: Precision measures the proportion of correctly identified positive instances among all predicted positives.
Recall: Recall measures the proportion of correctly identified positive instances among all actual positives.
F1-score: The F1-score balances precision and recall, serving as their harmonic mean. This metric is particularly useful for evaluating the trade-off between precision and recall in the DLA task.
FPS: FPS measures the number of frames processed by the model per second, providing an indication of the model’s inference speed and efficiency.
Train Dataset OCR MER TR DLA
[…]
Task-Specific 98.8 0.010 98.5 99.8 91.2 92.7 87.8
OCR+DLA 98.5 0.010 - - - - 88.9
OCR+MER+TR 99.1 0.008 98.9 99.9 92.3 94.6 -
@@ 5.3 Main Results
We use UnimerNet Wang et al. (2024a) for MER, StructEqTable Xia et al. (2024) for TR, DocLayout-YOLO Zhao et al. (2024c) for DLA, and Qwen-VL-OCR Bai et al. (2023) for OCR as baselines, as well as other widely used models for comparison. These baselines were selected for their strong performance and task relevance. The results show that DocFusion demonstrates competitive performance against other SOTA methods.
@@ 5.3.1 MER performance
We evaluated our model using the test subset of the UniMER-1M Wang et al. (2024a), with a focus on the Simple Printed Expression (SPE) and Complex Printed Expression (CPE) subsets, as DocFusion is specifically designed for processing printed documents. As shown in Table 2, DocFusion performs exceptionally well across multiple evaluation metrics, particularly in CSR and ExpRate. Notably, its ExpRate exceeds that of the second-ranked UniMERNet by 5.2%, demonstrating its superior reliability in real-world document parsing. The results presented here combine the performance of both SPE and CPE, with detailed separate results provided in the appendix.
@@ 5.3.2 TR performance
We constructed a benchmark dataset consisting of 2,500 table images extracted from LaTeX documents, including both simple and relatively complex tables, all of which were manually verified. To accommodate the model parameters and maximum sequence length, the LaTeX ground truth for the test set was limited to a maximum of 1,024 tokens. Using LatexNodes2Text, we extracted the content of each table cell to computed F1 scores (The detailed extraction method is presented in the appendix). As shown in Table 2, DocFusion performs excellently on this benchmark, with both F1 and CSR scores exceeding those of the second-ranked model by 1.6%, demonstrating its superior ability to handle both simple and complex table structures.
@@ 5.3.3 OCR performance
We separated 3,000 English image samples from the originally constructed dataset as the test set. As shown in Table 2, DocFusion demonstrates outstanding performance in both BLEU and EditDis, achieving more precise recognition of layout elements. This performance improvement is primarily attributed to DocFusion’s joint training on layout analysis and text recognition tasks, which enhances the model’s efficiency and effectiveness in handling complex document structures. These results further validate the effectiveness of the proposed training strategy, especially for document parsing tasks involving both text content and layout element recognition.
@@ 5.3.4 DLA performance
DocFusion generates layout element labels and coordinates by sequentially predicting tokens without relying on confidence scores. Since the commonly used Average Precision (AP) metric in object detection depends on confidence scores, it cannot be directly applied in this evaluation. To ensure a fair comparison with confidence-based models, we adopt an alternative evaluation method. For these models, we compute Precision, Recall, and F1-score at different thresholds and select the maximum F1-score across all thresholds as the final evaluation metric. As shown in Table 3, DocFusion may not achieve outstanding performance on the entire DocLaynet test set but performs well in the domain of scientific document detection. This could be attributed to its ability to generate bounding boxes with clean edges. In terms of processing speed, although DocFusion has more parameters than DETR, another Transformer-based model, it achieves faster processing due to the use of Flash-Attention. Compared to YOLO, DocFusion is slightly slower but does not require threshold tuning to achieve optimal performance, offering high performance without additional adjustments.
@@
CE 97.3 0.009 97.8 96.5 90.2 89.1 87.9
CE 99.1 0.007 98.7 99.8 92.1 92.5 88.2
Table 5: Ablation analysis of the improved objective function was conducted on the same dataset across four tasks: OCR, MER, TR, and DLA. CE represents training with the standard cross-entropy loss, while CE denotes training with the improved objective function.
@@ 5.4.1 OCR-Driven Enhancement of DLA
This section explores the impact of OCR on DLA performance. As shown in Table 4, the results in the DLA column from the first and second rows indicate that adding the OCR task improves DLA performance, with an F1 increase of up to 1.3%. This result demonstrates the effectiveness of using textual information in joint training. Compared to independent training that relies only on visual features, OCR significantly enhances the model’s robustness and generalization. For example, tables and mathematical expressions are layout elements with clear visual features, which the model can often recognize effectively. In contrast, text or titles have less distinctive visual features, making it challenging to predict their labels based on visual information alone. Textual cues play a crucial role in these cases. These findings confirm that OCR is essential for improving DLA performance. By providing complementary textual information, OCR strengthens the collaboration between visual and semantic features, resulting in better overall performance.
@@ 5.4.2 Collaboration of Recognition Tasks
In this section, we explore the collaboration among the recognition tasks OCR, TR, and MER. As shown in Table 4, the experimental results from the first and third rows demonstrate that joint training yields better performance compared to training each task individually. Specifically, OCR achieves a 0.3% improvement in BLEU score, MER sees increases of 0.4% in CDM and 0.1% in CSR, and TR benefits most significantly, with a 2.1% improvement in F1 score for cell parsing and a 2.0% increase in CSR. This collaboration enables the model to leverage shared information across tasks, enhancing individual task performance and improving overall document parsing capabilities. These results demonstrate that multi-task collaboration effectively enhances performance by leveraging shared information.
@@ 5.4.3 Results of improved objective function
In this section, we compared the original cross-entropy and the improved objective function in recognition and detection tasks. As shown in Table 5, the results demonstrate that the improved objective function led to significant performance gains across both task categories. In recognition tasks, the BLEU score in the OCR task saw an improvement of 1.8%. Additionally, the CDM metric in the MER task increased by 0.9%, while the F1 score in the TR task rose by 2.1%. Notably, for the CSR metric, which assesses LaTeX compilation success, the MER and TR tasks achieved gains of 3.3% and 3.8%, respectively, highlighting enhanced usability and correctness of the LaTeX outputs. For the detection task, the F1 score of the DLA task increased by 0.34%. This improvement can be attributed to the improved objective function, which alleviates the issue of coordinate token errors dominating the gradient. By addressing this imbalance, the objective function not only enhances the performance of recognition tasks but also improves the accuracy of predicting layout element categories in the detection task itself. These results collectively show that the improved objective function effectively addresses key challenges in loss minimization, ensuring that tasks such as DLA can operate within a generative framework. It avoids gradient dominance issues while achieving better task balance in a multi-task learning setup, demonstrating its robustness and versatility.
@@ 6 Conclusion
In this work, we introduced DocFusion, the first approach to integrate the four modules of a document parsing pipeline into a unified model by designing a objective function tailored to handle diverse data types across tasks. Our method achieved SOTA performance on multiple benchmarks. To enable downstream applications, we re-annotated the widely used DocLayNet dataset and constructed a large-scale formula-to-LaTeX dataset, applying a unified standardization process. Through detailed analysis, we observed that DocFusion, as a lightweight model with fewer parameters, effectively integrates multiple tasks into a single framework, demonstrating both efficiency and versatility in handling complex document parsing challenges. In the future, we aim to extend DocFusion to larger models and further improve dataset standardization to enhance its performance and applicability across broader tasks and domains.
@@
In this section, we discuss the limitations of the proposed model, DocFusion. While the model has demonstrated strong performance across multiple document layout analysis subtasks on specific datasets, its design is constrained by a parameter size of 289M and a maximum output length of 1024 tokens. These constraints may impact its ability to handle highly complex layouts or extremely long sequences, requiring further optimization for specific use cases. Additionally, DocFusion’s reliance on PDF screenshots in the LaTeX recognition task limits its generalization to handwritten or other non-standard formats. For the detection task, although the model achieves competitive accuracy, its processing speed poses challenges for real-time or high-throughput applications, indicating a need for further improvements in computational efficiency to meet broader application demands.
Appendix A Appendix
@@ A.1 DLA Dataset Reconstruction
In DocLaynet and other similar datasets, the annotation of mathematical formulas has certain limitations, as show in figure 5, the content of math expression and numbering are typically annotated within the same bounding box. This annotation approach introduces noise in subsequent Mathematical Expression Recognition (MER) tasks. To address this issue, we extracted formulas from arXiv LaTeX source files using regular expressions and assigned unique colors and bounding boxes to each element. Then, we employed a fuzzy matching algorithm to ensure annotation accuracy and eliminate overlaps. Finally, we trained a lightweight detection model and, combined with manual verification, re-annotated pages containing formulas. These improvements significantly enhance the dataset’s applicability to subsequent MER tasks.
A.2 MER and TR data standardization
@@
We chose to standardize the output format as LaTeX for two recognition tasks involving non-plain-text elements. For MER, converting to LaTeX was essential as it provides a precise representation of mathematical formulas. For TR, in addition to ensuring format consistency, converting to LaTeX also allows for the restoration of the original content through compilation, and enables the extraction of cell elements using tools such as LatexNodes2Text, thus enhancing processing flexibility. We used regular expressions to extract relevant content from the LaTeX source files of research papers. However, due to variations in author writing styles, the same formula or table may appear in multiple forms, increasing the complexity of training. As show in table 6 , we analyzed these different representations, standardized them to eliminate ambiguities and ensured consistency. To verify the accuracy of the standardized LaTeX code, we re-rendered it into images, creating a high-quality dataset that aligns with the actual input-output content.

### v2

@@ 4.5.1 MER performance
We use the open-source UniMER-1M Wang et al. (2024b) as the evaluation dataset to assess the performance on MER. Since DocFusion is specifically designed for processing printed documents, the primary evaluation focuses on the Simple Printed Expression (SPE) and Complex Printed Expression (CPE) subsets of UniMER-1M. As shown in Table 2, DocFusion performs exceptionally well across multiple evaluation metrics, particularly in CSR and ExpRate. Notably, its ExpRate surpasses the second-ranked UniMERNet by 5.2%, demonstrating superior reliability in real-world document parsing. The results presented here merge the performance of both SPE and CPE, with detailed separate results and handwritten expressions provided in B.1.
@@ 4.5.2 TR performance
We selected DocGenome Xia et al. (2024) as the evaluation dataset because it offers a comprehensive collection of 500K scientific documents across various disciplines, covering a wide range of document-oriented tasks. From this dataset, we extracted 3,000 LaTeX-based table samples as the test set. Using LatexNodes2Text, we extracted the content of each table cell to compute F1 scores. As shown in Table 2, DocFusion excels on this benchmark, achieving F1 scores that surpass those of the second-ranked model by 1.6%, while having less than one-third of its parameter count. Note: In this work, in order to maintain consistency with MER and explore multi-task collaboration, we also chose Latex as the output format for our TR task. However, in the past, Latex was not mainstream in Table-to-Sequence tasks, so there are fewer models available for comparison. To provide more comprehensive reference information, we have included the F1 scores of other models that output in HTML in the appendix B.2.
@@ 4.5.3 OCR performance
As mentioned in 4.1, DocLayNet Pfitzmann et al. (2022) supports not only DLA but also OCR evaluation. We selected 3,000 English image samples from the dataset as the test set. As shown in Table 2, DocFusion achieves exceptional performance in both BLEU and EditDis. This outstanding result is primarily attributed to DocFusion’s joint training on three recognition tasks, which enhances its efficiency and effectiveness in handling complex document structures.
@@ 4.5.4 DLA performance
We use the test set from DocLayNetPfitzmann et al. (2022) to evaluate the DLA task. In terms of FPS, while DocFusion exhibits a slight disadvantage in processing speed, it offers an out-of-the-box solution that eliminates the need for hyperparameter tuning in practical applications. This enables the model to achieve optimal performance directly, without requiring further adjustments, thereby compensating for its lower speed.
Regarding accuracy, DocFusion generates layout element labels and coordinates by sequentially predicting tokens without relying on confidence scores. Given that the commonly used Average Precision (AP) metric in object detection is based on confidence scores, it is not directly applicable in this evaluation. To ensure a fair comparison with confidence-based models, we adopt an alternative evaluation methodology. Specifically, for these models, we compute Precision, Recall, and F1-score at various thresholds and select the maximum F1-score across all thresholds as the final evaluation metric. As shown in Table 3, DocFusion demonstrates strong performance in the domain of scientific document detection.
@@
CE 96.5 2.3 97.8 96.5 90.2 89.1 87.9
GK-CEL 97.4 1.8 98.7 99.8 92.1 92.5 88.4
Table 5: Ablation analysis of Gaussian-Kernel Cross-Entropy Loss was conducted on the same dataset across four tasks: OCR, MER, TR, and DLA. CE represents training with the standard cross-entropy loss, while GK-CEL denotes training with Gaussian-Kernel Cross-Entropy Loss.
@@ 4.6.1 OCR-Driven Enhancement of DLA
This section explores the impact of OCR on DLA performance. As shown in Table 4, the results in the DLA column from the first and second rows indicate that adding the OCR task improves DLA performance, with an F1 increase of up to 1.3%. This result demonstrates the effectiveness of using textual information in joint training. Compared to independent training that relies only on visual features, OCR significantly enhances the model’s robustness. For example, tables and mathematical expressions have distinct visual features, which can often be effectively recognized by DLA models. In contrast, text or titles have less distinctive visual features, making it challenging to predict their labels based on visual information alone. By providing complementary textual information, OCR strengthens the collaboration between visual and semantic features, resulting in better overall performance.
@@ 4.6.2 Collaboration of Recognition Tasks
In this section, we explore the collaboration among the recognition tasks OCR, TR, and MER. As shown in Table 4, the experimental results from the first and third rows demonstrate that joint training yields better performance compared to training each task individually. Specifically, OCR achieves a 0.4% improvement in BLEU score, MER sees increases of 0.4% in CDM and 0.1% in CSR, and TR benefits most significantly, with a 1.2% improvement in F1 score for cell parsing and a 2.1% increase in CSR. This collaboration enables the model to leverage shared information across tasks, enhancing individual task performance and improving overall document parsing capabilities. These results demonstrate that multi-task collaboration effectively enhances performance by leveraging shared information.
@@ 4.6.3 Results of improved objective function
In this section, we compared the original cross-entropy and Gaussian-Kernel Cross-Entropy Loss (Gk-CEL) in recognition and detection tasks. As shown in Table 5, the results demonstrate that Gk-CEL led to significant performance gains across both task categories. In recognition tasks, the BLEU score in the OCR task saw an improvement of 0.9%. Additionally, the CDM metric in the MER task increased by 0.9%, while the F1 score in the TR task rose by 2.1%. Notably, for the CSR metric, which assesses LaTeX compilation success, the MER and TR tasks achieved gains of 3.4% and 3.8%, respectively, highlighting enhanced usability and correctness of the LaTeX outputs. For the detection task, the F1 score of the DLA task increased by 0.5%. This improvement can be attributed to Gk-CEL, which alleviates the issue of coordinate token errors dominating the gradient. By addressing this imbalance, the objective function not only enhances the performance of recognition tasks but also improves the accuracy of predicting layout element categories in the detection task itself. These results collectively show that Gk-CEL effectively addresses key challenges in loss minimization, ensuring that tasks such as DLA can operate within a generative framework. It avoids gradient dominance issues while achieving better task balance in a multi-task learning setup.
@@ 5 Conclusion
In this work, we introduced DocFusion, the first approach to integrate the four modules of a document parsing pipeline into a unified model by designing Gaussian-Kernel Cross-Entropy Loss tailored to handle diverse data types across tasks. Our method achieved SOTA performance on multiple benchmarks. To enable downstream applications, we re-annotated the widely used DocLayNet dataset and constructed a large-scale formula-to-LaTeX dataset, applying a unified standardization process. Through detailed analysis, we observed that DocFusion, as a lightweight model, effectively integrates multiple tasks into a single framework, demonstrating both efficiency and versatility in handling complex document parsing challenges. In the future, we aim to extend DocFusion to larger models and further improve dataset standardization to enhance its performance and applicability across broader tasks and domains.
@@
While this study primarily focuses on three recognition tasks using standard PDF screenshots, we have enhanced the model’s generalization capabilities by incorporating handwritten mathematical expressions. However, the model still has limitations in handling handwritten or other non-standard table formats. For the detection task, although the model demonstrates competitive performance in both accuracy and usability, its processing speed presents challenges for real-time or high-throughput applications. This highlights the need for further optimization in computational efficiency to better meet diverse application demands.
Acknowledgements
TThe authors wish to thank the anonymous reviewers for their helpful comments. This work was funded by National Natural Science Foundation of China (No.62476061)
Appendix A Details of Datasets
@@ A.1 DLA Dataset Reconstruction
In DocLaynet and other similar datasets, the annotation of mathematical formulas has certain limitations, as show in figure 5, the content of math expression and numbering are typically annotated within the same bounding box. This annotation approach introduces noise in subsequent Mathematical Expression Recognition (MER) tasks.
To address this issue, we extracted formulas from arXiv LaTeX source files using regular expressions and assigned unique colors and bounding boxes to each element. Then, we employed a fuzzy matching algorithm to ensure annotation accuracy and eliminate overlaps. Finally, we trained a lightweight detection model and, combined with manual verification, re-annotated pages containing formulas. These improvements significantly enhance the dataset’s applicability to subsequent MER tasks.
A.2 MER and TR Dataset
MER Dataset. The UniMER-1M Wang et al. (2024b) has significantly advanced MER research but contains many redundant spaces in LaTeX code. Although some spaces are syntactically necessary, most are unnecessary, increasing output length and computational overhead. To address this, we constructed a new dataset by extracting content from LaTeX files, normalizing style variations and verifying accuracy through re-rendering. Models trained on our dataset produce LaTeX code that is approximately 34.2% shorter for complex expressions and 37.5% shorter for simple expressions on the UniMER-1M test set, demonstrating improved efficiency.
A.3 Latex-based data standardization
@@
We chose to standardize the output format as LaTeX for two recognition tasks involving non-plain-text elements. For MER, converting to LaTeX was essential as it provides a precise representation of mathematical formulas. For TR, in addition to ensuring format consistency, converting to LaTeX also allows for the restoration of the original content through compilation, and enables the extraction of cell elements using tools such as LatexNodes2Text, thus enhancing processing flexibility.
We used regular expressions to extract relevant content from the LaTeX source files of research papers. However, due to variations in author writing styles, the same formula or table may appear in multiple forms, increasing the complexity of training. As show in table 6 , we analyzed these different representations, standardized them to eliminate ambiguities and ensured consistency. To verify the accuracy of the standardized LaTeX code, we re-rendered it into images, creating a high-quality dataset that aligns with the actual input-output content.
Model size SPE CPE HWE
[…]
Pix2tex (2022) - 92.1 59.0 99.8 45.2 7.2 88.1 24.7 8.1 16.3
Texify (2023) 312M 98.7 89.8 99.8 69.8 35.6 94.3 49.9 21.3 25.8
GOT (2024) 580M 95.0 82.7 98.6 73.3 36.4 96.4 31.2 17.7 10.2
UniMERNet (2024b) 325M 99.7 95.6 99.9 97.6 77.4 99.2 94.7 65.3 98.1
DocFusion(Ours) 289M 99.7 97.3 99.9 96.9 88.1 99.5 94.1 72.1 99.3
Table 7: Supplementary details of MER. SPE refers to simple printed mathematical expressions, CPE refers to complex printed mathematical expressions, and HWE refers to handwritten mathematical expressions.
Appendix B Other supplementary experiments
B.1 Details of MER Performance
we provide a detailed presentation of the main experimental results for MER, showing the performance of the relevant models on simple, complex, and non-standard handwritten mathematical expressions. For specifics, please refer to Table 7.
B.2 Other Table-to-Sequence Method
Methods F1 CSR
surya 37.4 -
ppstructure_table 78.1 -
Deepdoctection 53.7 -
RapidTable 87.9 -
MixTex 46.2 27.4
GOT 86.9 81.6
StructEqTable 90.6 93.2
DocFusion 92.1 92.5
Table 8: Due to differences in the method of extracting cell contents, the fairness of the experiment cannot be guaranteed, therefore, it is provided for reference only.
This study aims to explore multi-task collaboration, and therefore, the TR task also adopts Latex as the output format to maintain consistency with MER. However, Latex has not been the mainstream approach for TR tasks in recent times, resulting in a limited number of TR models available for comparison in the main experiment. To address this limitation, we incorporated other methods based on HTML as the output format. However, due to differences in sequence extraction methods, ensuring a fair comparison is challenging. Therefore, we have included the supplementary experimental results in the appendix for reference.
Appendix C Other optimization methods
The challenge of this experiment lies in effectively optimizing continuous coordinate-type data within a discrete generative framework. Since there are inherent errors in coordinate annotations, these errors are further amplified when training the generative framework using cross-entropy loss, especially when the framework performs multiple tasks, which exacerbates the issue. To address this problem, in addition to the Gaussian-Kernel Cross-Entropy Loss introduced in the main text, we employed several other optimization strategies, including the basic adjustments of data ratios or loss weights, as well as using soft-argmax to continuously map discrete coordinate tokens.
C.1 Hyperparameters Adjustment Strategies
The root cause of the training difficulty lies in the fact that the discrete coordinate tokens do not effectively dominate the loss during training, leading to poor gradient propagation and inefficient parameter updates. To address this, one possible solution is to adjust the data ratios or the loss weights across different task types. However, while this approach can improve training stability to some extent, it is overly engineering-driven and does not fundamentally solve the underlying issue of inadequate gradient flow caused by the discrete nature of the coordinate tokens.
Model size OCR MER TR DLA
[…]
DocFusion-base 289 97.4 1.8 98.7 99.8 92.1 92.5 88.4
DocFusion-large 738 97.2 1.9 99.1 99.8 92.4 92.5 89.1
Table 9: Ablation analysis of Gaussian-Kernel Cross-Entropy Loss was conducted on the same dataset across four tasks: OCR, MER, TR, and DLA. CE represents training with the standard cross-entropy loss, while GK-CEL denotes training with Gaussian-Kernel Cross-Entropy Loss.
C.2 Soft-argmax Strategies
Figure 6: Soft-argmax Loss
The core issue lies in the fact that while multi-task frameworks need to be discrete, coordinates are inherently continuous. A natural solution to this problem is to "smooth" the coordinate loss, effectively making it continuous. This approach offers an intuitive way to handle the challenge, and we primarily use the soft-argmax technique to obtain the position coordinates while maintaining the gradient flow, followed by the computation of the loss via Mean Squared Error (MSE).
However, the difficulty arises during multi-task training: after calculating the MSE, we need to ensure that it remains within the same range as other cross-entropy (CE) losses. The challenge here is to maintain balance and prevent the MSE loss from overwhelming the CE losses. Moreover, if the hyperparameters of the soft-argmax are not set appropriately, it can easily lead to gradient explosion during training, further complicating the optimization process.
Although this method aims to address the issue at its core by making the coordinate loss continuous, it still relies heavily on the correct setting of hyperparameters. Furthermore, it presents generalization issues when applied to different tasks or datasets. In comparison, the Gaussian-Kernel Cross-Entropy Loss (GK-CEL) offers a more robust solution, as it reduces the dependency on hyperparameters while improving generalization performance.
C.3 Model Size Analysis
While model performance generally benefits from increased parameter size, the advantages can diminish in recognition-oriented tasks due to the limited gains in accuracy relative to the added computational cost. In the early stages of this work, we trained a larger version of our model with 738M parameters. Although it achieved slightly better performance on certain tasks—such as a modest improvement on the DLA benchmark—the gains were not substantial enough to justify the significantly higher inference cost, especially given the autoregressive nature of decoding.
As our primary goal is to demonstrate the feasibility of a lightweight and unified model, we chose to focus on the 289M version of DocFusion in this paper. As shown in Table9, this smaller model already delivers strong results across tasks, including 97.4 BLEU on OCR and 99.8 CSR on MER. We believe this better reflects the practical trade-off between efficiency and performance. Results from the larger variant will be included in a future version to facilitate further exploration and provide a reference for the research community.
Figure 7: DLA Effect Presentation
Figure 8: MER Effect Presentation
Figure 9: TR Effect Presentation